OpenStudy (anonymous):

Eliminate the parameter and obtain the rectangular equation if x=7t-5 y=5t+5

jimthompson5910 (jim_thompson5910):

Solve x=7t-5 for t x=7t-5 x+5=7t 7t = x+5 t = (x+5)/7

jimthompson5910 (jim_thompson5910):

you then plug this into the second equation y=5t+5 y=5*(x+5)/7+5 I'll let you simplify

OpenStudy (anonymous):

Is the answer y=(5/7)x + 60/7 ?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

you're welcome